KrazyKat02:

Avery owns stocks in Ever Rise Corporation. The stock price increased by 120% over the past week. If the stock price was p last week, write two equivalent expressions that represent the current stock price and explain what their equivalence means. Substitute one or more variables into the expressions to prove them equivalent.

LumineXgirl:

P(1.2) and p+ p(0.2) are equivalent expressions because they both equal to p(1.2). P=2 Expression 1: P(1.2)=2(1.2)=2.4 Expression 2: P+p(0.2)= 2 + 2(0.2)= 2.4 2.4=2.4

KrazyKat02:

Thanks!

LumineXgirl:

np